### v1

# Dual-Domain Homogeneous Fusion with Cross-Modal Mamba and Progressive Decoder for 3D Object Detection

Abstract
Fusing LiDAR point cloud features and image features in a homogeneous BEV space has been widely adopted for 3D object detection in autonomous driving. However, such methods are limited by the excessive compression of multi-modal features. While some works explore feature fusion in dense voxel spaces, they suffer from high computational costs and inefficiencies in query generation. To address these limitations, we propose a Dual-Domain Homogeneous Fusion network (DDHFusion), which leverages the complementary advantages of both BEV and voxel domains while mitigating their respective drawbacks. Specifically, we first transform image features into BEV and sparse voxel spaces using LSS and our proposed semantic-aware feature sampling module which can significantly reduces computational overhead by filtering unimportant voxels. For feature encoding, we design two networks for BEV and voxel feature fusion, incorporating novel cross-modal voxel and BEV Mamba blocks to resolve feature misalignment and enable efficient yet comprehensive scene perception. The output voxel features are injected into the BEV space to compensate for the loss of 3D details caused by height compression. For feature decoding, a progressive query generation module is implemented in the BEV domain to alleviate false negatives during query selection caused by feature compression and small object sizes. Finally, a progressive decoder can sequentially aggregate not only context-rich BEV features but also geometry-aware voxel features, ensuring more precise confidence prediction and bounding box regression. On the NuScenes dataset, DDHfusion achieves state-of-the-art performance, and further experiments demonstrate its superiority over other homogeneous fusion methods.
Index Terms:
3D object detection, autonomous driving, multi-modal, homogeneous fusion, Mamba.
I Introduction
Autonomous driving has emerged as a highly active research field, advancing rapidly with the integration of deep learning models. Among its critical technologies, 3D object detection plays a pivotal role in enabling safe and efficient driving planning. In current autonomous systems, LiDAR and cameras serve as essential sensors for environmental perception. LiDAR utilizes laser echo technology to capture precise 3D environmental information in the form of point clouds, while cameras generate images rich in texture information by capturing visible light. By harnessing the complementary strengths of LiDAR and cameras, fusion-based object detectors have consistently outperformed their single-sensor counterparts in terms of accuracy and robustness. However, multi-sensor feature fusion is non-trivial due to the inherent heterogeneity between camera and LiDAR data. Many existing works [1, 2, 3, 4, 5, 6] attempt to enrich raw point clouds or voxels with image features. However, this approach suffers from semantic information loss during camera-to-LiDAR projection, as only a small fraction of image features are preserved, leading to degraded detection performance in regions with fewer points. Recent studies propose homogeneous fusion strategies [7, 8, 9, 10, 11, 12] that actively transform image features into a shared representation space with LiDAR features. These methods leverage symmetric network architectures to fully exploit multi-modal complementarity while mitigating performance degradation caused by single-sensor failures.
Figure 1: Overview of homogeneous fusion methods. The view transformation in Fig. 1(a) and Fig. 1(b) causes the loss of modality details. The dense voxel representation in Fig. 1(c) causes computational burden. Our method in Fig. 1(d) combines the advantages of sparse voxel and BEV domains.
A popular method [7, 8, 9], as depicted in Fig. 1(a), employs Lift-Splat-Shoot (LSS) to project image features onto a unified BEV plane for concatenation with LiDAR features. This approach is highly efficient as mature techniques from 2D detection tasks can be easily adopted for BEV feature fusion and query generation. For the latter, the sample center-based label assignment and Top-K selection strategies are widely applied. However, it also has two limitations: 1) Information Compression: The compression of geometric and texture information along the height dimension results in blurred BEV features, which further leads to missed detections and reduced regression accuracy. As shown in Fig. 1(b), some works [10, 11] introduce multi-view encoder-decoder structures to address this. However, the additional transformation of LiDAR features back to the frontal view significantly disrupts 3D spatial relationships, still incurring semantic ambiguity. 2) Feature Misalignment: Inaccurate depth estimation in LSS or calibration errors between sensors may lead to spatial misalignment of multi-modal BEV features. Extra alignment techniques are required before feeding the features into the BEV backbone.
As shown in Fig. 1(c), fusion in a unified voxel space has been explored to retain 3D geometric details and leverage higher-dimensional feature interactions for enhanced spatial perception while avoiding information compression. While theoretically superior to BEV-based fusion, this paradigm faces several challenges: 1) Computational Burden: Fusing features represented as dense voxels is highly memory-intensive and computationally expensive, making it difficult to deploy in large-scale driving scenarios. 2) Query Generation Difficulty: According to [13], compared to the BEV domain, the more severe foreground-background imbalance and higher dimensionality in the voxel domain complicate label assignment and input-dependent query generation [14]. [12] adopts randomly initialized 3D learnable queries combined with local 3D deformable attention in decoders, which slows down training convergence and limits the model’s ability to localize potential instances. 3) Feature Misalignment: The feature misalignment issue becomes even more challenging in the voxel domain. Computational constraints force these methods to rely on 3D convolutions for local feature fusion within limited receptive fields, which hinders feature alignment over larger regions.
In this work, we propose a novel detection framework—the Dual-Domain Homogeneous Fusion network (DDHFusion), which overcomes the limitations of both BEV and voxel fusion domains while leveraging their respective strengths. As shown in Fig. 1(d), the BEV domain is responsible for query generation, while the voxel domain provides geometry-aware features. For clarity, the proposed method is divided into three steps: 1) Feature Transformation: First, we transform image features into sparse voxel and BEV spaces using the semantic-aware feature sampling (SAFS) and LSS modules. Compared to [12], the sparse voxel representation significantly reduces computational overhead. 2) Feature Encoding: During encoding, these multi-representation image features interact with LiDAR ones through homogeneous voxel and BEV fusion (HVF and HBF) networks. Within these networks, we introduce cross-modal voxel and BEV mamba blocks (CV-Mamba and CB-Mamba) to resolve feature misalignment and enable efficient global scene perception. 3) Feature Decoding: For decoding, we propose a progressive query generation (PQG) module and a progressive decoder (PD). The former activates BEV features with high-response queries, generating potential hard queries in the second stage to mitigate false negatives caused by feature compression and small object sizes. The latter enhances the conventional BEV decoder by integrating fine-grained 3D structural features around targets with our proposed multi-modal voxel feature mixing (MMVFM) module.
As a summary, the main contributions are as follows:
  1. 1.
    We introduce DDHFusion for 3D object detection that achieves state-of-the-art performance on the NuScenes dataset, outperforming other homogeneous fusion approaches while maintaining efficient inference.
  2. 2.
    To the best of our knowledge, we are the first to extend Mamba into multi-modal 3D detection. The proposed CV-Mamba and CB-Mamba blocks effectively facilitate feature fusion across global voxel or BEV domains.
  3. 3.
    We propose the PQG and PQ modules for feature decoding. The former generates easy and difficult queries in stages to minimize false negatives, while the latter adaptively integrates dual-domain features to enhance the accuracy of each query’s regression.
II related work
In this section, we overview homogeneous fusion methods in the BEV and voxel domains, followed by recent advancements in state space models.
II-A Homogeneous Fusion in BEV Domain
Fusion in a shared BEV space is a prominent approach for 3D object detection in autonomous driving. A critical step in this process is the transformation of image features into the BEV space. Early work [15] proposed an auto-calibration method: image features are projected onto predefined dense 3D voxel centers, adjusted by regressing offsets, and assigned to corresponding voxels. These features are then compressed along the height dimension to generate BEV features. Building on this, [16] introduced range attention to refine the distribution of intermediate voxel features along the range dimension. Recently, LSS [17] has become widely adopted for view transformation. It projects features into a 3D frustum based on depth estimation and then onto the BEV place with pooling operation. However, inaccuracies in depth estimation often lead to misalignment between LiDAR and image BEV views. To address this, [18] leverages contrastive learning for feature alignment, while [19] and [20] utilize semantic-guided optical flow estimation or mutual deformable attention for explicit alignment. Additionally, [21] and [20] enhance depth estimation accuracy through edge-aware LiDAR depth maps or local alignment operations.
Unlike the local fusion operations in the aforementioned methods, our approach establishes global relationships both within and across modalities, enabling effective feature alignment and more comprehensive scene perception.
II-B Homogeneous Fusion in Voxel Domain
Compared to BEV homogeneous fusion, transforming image features into the voxel domain and fusing them with LiDAR voxels avoids the information loss caused by height compression. Additionally, introducing multi-modal voxel features during the decoding process leads to better regression accuracy. [22] and [23] unproject image features from each pixel onto virtual points with depth completion and then convert them into voxels, applying voxel pooling to fuse fine-grained multi-modal features. [24] generates virtual points using nearest neighbor matching around the 2D instance mask on the image. [25] improves this approach with k-nearest neighbor matching and introduces gated modality-aware convolution to fuse semantic and geometric features of both camera and LiDAR at different granularities. While these virtual-point-based methods largely address the sparsity issue of LiDAR point clouds, they require additional depth completion or high-resolution instance segmentation networks. It leads to significant training costs, and the labels required for these tasks are often difficult to obtain. In contrast, [12] bypasses the need for extra networks by adopting an LSS-like algorithm. It first assigns image features to predefined dense voxels and then weights them with category-based depth maps. [26] eliminates the depth-weighting step and introduces additional gating mechanisms to enable flexible fusion. However, dense voxel representations incur significant computational overhead. In recent work, [27] only utilize element-wise fusion in the intermediate voxel space during LSS, with subsequent feature propagation occurring in the BEV domain.
In our work, we adopt sparse voxel representation to conserve computational resources. Additionally, our Mamba-based network, HVF, facilitates multi-granularity global feature fusion.
II-C State Space Model
State Space Models (SSMs) [28, 29, 30] describe dynamic systems by representing the system’s internal state at each moment. With advancements in deep learning, SSMs have been integrated into neural networks to achieve more efficient and flexible inference. [31] effectively captures long-term dependencies in sequences by introducing higher-order polynomial projection operators, compressing historical information into a low-dimensional state space. S4 [28], a structured state space model based on [29], enhances computational efficiency and modeling capacity through structured state transition matrices. Building on S4, [32] introduces input-specific state transition mechanisms, significantly boosting both expressiveness and efficiency. Due to its linear complexity, Mamba has been widely adopted in computer vision. For instance, [33] unfolds images into 1D sequences and proposes a bidirectional Mamba module for comprehensive feature learning. [34] extends this approach to four-directional scanning, uncovering richer spatial relationships in images and providing a theoretical analysis of Mamba’s linear attention. Mamba has also been applied to 3D tasks such as point cloud classification [35], 3D object detection [36, 37, 38], and semantic completion [39]. For example, [36] employs a group-free voxel Mamba module to expand the receptive field across the entire scene, while [39] integrates local adaptive reordering into the Mamba module to enhance local information extraction.
Different from these single-modality architectures for 3D tasks, we extend Mamba to multi-modal ones, further unlocking its potential.
III Method
The pipeline of DDHFusion is shown in Fig. 2. First, multi-view images and LiDAR point clouds are passed through their respective networks for feature extraction. These features are then transformed into sparse voxels $V_{I}$, $V_{L}$ and BEV representations $B_{I}$,$B_{L}$. Then, two homogeneous fusion networks are introduced to perform feature fusion within their respective domains. The outputs $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$ are passed into the BEV fusion network after sparse height compression [40]. After processing through the BEV backbone, the resulting BEV features are used for progressive query generation. These queries are then input into the progressive decoder including both BEV and voxel decoders for classification and box regression. In this section, we provide a detailed introduction to DDHFusion from the following aspects: semantic-aware feature sampling, voxel homogeneous network, BEV homogeneous network, progressive query generation, and progressive decoder.
III-A Semantic-Aware Feature Sampling
Describing the entire driving scene using dense voxels is highly redundant. First, a significant portion of voxels corresponds to empty space, where no objects are present. Second, in autonomous driving scenarios, foreground objects typically occupy only a small fraction of the overall scene. This redundancy not only consumes substantial computational resources but also diverts the model’s information extraction capacity to large, irrelevant areas, thereby hindering feature learning.
Drawing inspiration from [41], we propose the SAFS module depicted in Fig. 3 to selectively generate important image voxels. First, the image features $F_{I}$ from the image backbone are passed through a convolution block to generate the depth map. Similar to LSS, these depths are represented by discrete bins. Another branch predicts the semantic segmentation mask. During training, we project points within ground truth boxes onto the image plane to generate segmentation labels for sparse supervision.
Figure 2: Overview of DDHFusion. It begins by extracting features from multi-view images and LiDAR point cloud and transforms both into the voxel and BEV domains. These features are then passed to two Mamba-based homogeneous fusion networks, which perform feature alignment and global perception. The resulting high-quality fused BEV feature $B_{out}$ is then fed into the progressive query generation module, which leverages the spatial relationship of easy queries to stimulate the generation of hard queries. Finally, all queries are passed into the progressive decoder, which abstracts dual-domain features around instances for accurate classification and box regression.
Figure 3: Details of semantic-aware feature sampling (SAFS). We filter unimportant voxels by depth and semantic scores and transform the image feature to the sparse voxel domain.
We divide the entire 3D space into voxels $V_{c}$ with number $H_{I}\times W_{I}\times Z_{I}$. According to [12], we set the height range $Z_{I}$ twice as that of LiDAR voxels $Z_{L}$ to obtain more texture details from images. Then, we project voxel centers onto the image plane and collect depth scores $V_{d}$ and semantic score $V_{s}$ using bilinear interpolation. The voxels with both scores above certain thresholds are selected:
$V_{c}=\{{v_{ci}}\mid({v_{si}}>s)\land({v_{di}}>d)|$
$\quad{v_{ci}}\in{V_{c}},\,{v_{di}}\in{V_{d}},\,{v_{si}}\in{V_{s}},\,i\in[0,{H_{I}}{W_{I}}{Z_{I}}-1]\}$ (1)
At the beginning of training or in some crowd scenarios, the number of $V_{c}^{{}^{\prime}}$ may be too large. Therefore, we set an upper limit $N$, and if the number exceeds it, the farthest point sampling is applied to retain only $N$ voxels. Finally, we assign image features to $V_{c}^{{}^{\prime}}$ and multiply them by the corresponding depth scores to obtain voxel features $V_{I}$.
III-B Homogeneous Voxel Fusion Network
In HVF, we employ a 3D U-Net architecture to fuse LiDAR voxels $V_{L}$ and image voxels $V_{I}$. As illustrated in Fig. 4, the network consists of two parallel branches that process voxel features separately, with feature fusion performed at each scale. Due to the uncertainty of depth estimation and calibration error, $V_{L}$ and $V_{I}$ are often misaligned spatially. Previous works [12, 22, 25] only fuse voxel features by concatenation or local gating mechanisms within the same position or a $3\times 3\times 3$ neighborhood. However, they fail to resolve the misalignment problem and limit the network’s ability to comprehensively perceive multi-modal information. Inspired by the recent success of Mamba in LiDAR-based 3D object detection [36, 37], we propose intra-modal and cross-modal voxel Mamba (IV-Mamba and CV-Mamba) modules for feature fusion at each scale. In IV-Mamba, we rearrange the voxels into a 1D sequence based on the 3D Hilbert curve according to [36]. The discrete SSM model is then used to process the entire sequence:
$\left\{\begin{array}[]{l}h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i}\\ y_{i-1}=\overline{C}h_{i}+x_{i}\end{array}\right.$ (2)
Figure 4: Pipeline of Mamba-based Voxel Fusion. We utilize the bidirectional Mamba block, as illustrated in the right figure, to process both single-modal and multi-modal voxel features. Within the 3D U-Net architecture, the merge-and-split operation seamlessly switches between intra-modal and cross-modal Mamba blocks, facilitating global alignment and fusion of multi-modal information.
Figure 5: Details of Intra-Modal and Cross-Modal BEV Mamba. We apply a 4-directional cross-scan to unfold the BEV feature into 1D sequences, which helps construct a more comprehensive spatial relationship. In the cross-modal Mamba block, the parameters for the SSMs of each modality are jointly computed from the Image and LiDAR BEV features, while $Z_{I}$ and $Z_{L}$ are used to adaptively modulate the weights of their respective modalities.
Here, $\overline{A}$ and $\overline{B}$ are derived from the parameters of the continuous SSM using the zero-order hold formula:
[…]
Where $\bigtriangleup$ denotes the time-scale parameter and $A$ denotes the learnable matrix. $B$, $C$, and $\bigtriangleup$ are inferred from the input sequence $x_{i}$ through linear layers, ensuring the input-specific of the SSM model. To ensure complete perception, both forward and backward SSMs are implemented. Their output sequences are added together and modulated by $Y$ to produce the final results.
In CV-Mamba, we directly combine multi-modal voxels in the unified 3D space. It’s important to note that the height range of image voxels $Z_{I}$ is twice that of LiDAR voxels $Z_{L}$. Therefore, we need to multiple the z indices of LiDAR voxels by two before combination. Similar to IV-mamba, these voxels are unfolded into 1D sequences and fed into the SSM. While some voxels may occupy the same position, we treat them as distinct elements in the sequence rather than merging them as in [25]. According to the theoretical proof in [34], the forward-backward Mamba is a special version of a global transformer, enabling unaligned features to match their corresponding ones in the other modality. Then, the output voxels with different modalities are separated and sent back to their perspective branches, where they are downsampled by sparse 3D convolution. Notably, in most existing works, fusing features with different distributions[23, 25, 22, 42] requires operations such as nearest-neighbor search or grouping, while our designed elegant global merge-and-split operation makes the fusion process more efficient and straightforward. Alternating between IV-Mamba and CV-Mamba effectively bridges the modality gap and alleviates the negative impact of deteriorated data from each modality on the fused features.
Finally, Inverse sparse convolutions are used to upsample voxels features back to the original scale. The output features $V_{I}$ and $V_{L}$ are not directly combined to avoid interfere brought by distinctive voxel distributions. Instead, they are separately fed into the homogeneous BEV fusion network and the multimodal feature mixing module in the decoder.
III-C Homogeneous BEV Fusion Network
In Homogeneous BEV Fusion Network, the features from the homogeneous voxel network, $V_{I}^{’}$ and $V_{L}^{’}$, are transformed into BEV grids through sparse height compression [43] — specifically, max-pooling is applied in voxel features belonging to the same BEV pixel, yielding $B_{I}^{{}^{\prime}}$ and $B_{L}^{{}^{\prime}}$. They are then concatenated with $B_{I}$ and $B_{L}$ to inject more 3D semantic and structure details into BEV space. It is important to note that during SAFS, some potential foreground voxels may be filtered out due to low semantic or depth scores. The LSS compensates for this by retaining image information of the entire environment. In addition, both SAFS and LSS share the same depth map. Since LSS has not undergone feature filtering, it preserves all the backpropagated gradients, which helps accelerate the convergence of depth estimation during training.
As shown in Fig. 4, we also applied the Mamba-based fusion paradigm, which includes both intra-modal and cross-modal BEV Mamba (IB-Mamba and CB-Mamba) modules to generate high-quality BEV features. First the concatenated features into the IB-Mamba, enabling the network to globally perceive modality-specific information. Following the SS2D block proposed in [34], we apply four-directional cross scanning to unfold the image, which helps construct a comprehensive spatial relationship. Then, as described in Equations 2 and 3, the SSM is performed simultaneously across all four directions. After passing through LayerNorm, the outputs are summed and modulated by $Y$ for final results.
In CB-Mamba, to further align multi-modal features in the dense BEV space and adaptively fuse them, we extend the SS2D operation originally designed for single-modal tasks to cross-modal fusion. As shown in Fig. 4, the parameters for the SSM computation are generated from the concatenated tensor, guiding the global rectification for both image and LiDAR features. We describe the SSM parameterization using the following formulas.
$T=Conv2d(SiLU(BatchNorm(Conv2d(F_{comb}))))$ (4)
$[B_{I}^{1-4},C_{I}^{1-4},\bigtriangleup_{I}^{1-4},B_{L}^{1-4},C_{L}^{1-4},\bigtriangleup_{L}^{1-4}]=Split(T)$ (5)
Where, the superscripts $1-4$ denote the parameters employed for the four-directional scanning in SS2D. Furthermore, the modulation parameters $Y_{I}$ and $Y_{L}$ are also generated by concatenating features and are used to reweight multi-modal features:
$Y_{I}=SiLU(Conv2d(T))$ (6)
$Y_{L}=I-Y_{I}$ (7)
It enables the network to autonomously favor the more reliable modality, thereby improves adaptability in complex scenarios.
After CB-Mamba, the output BEV tensor is fed into the BEV backbone for further feature propagation, yielding $B_{out}$, which is then utilized as the input for query generation.
Figure 6: Details of Hard Instance Activation. First, self-attention among easy queries is used to extract spatial relationships between objects. Then, query-to-BEV attention is applied to activate latent hard instances within the BEV features.
III-D Progressive Query Generation
In DDHFusion, the BEV branch is used for efficient generation of target candidates. While many previous works have attempted to achieve this in the range view[44, 45] or sparse 3D space[46, 47], the former is impacted by object occlusions, and the latter often relies on center-voting strategies, which are sensitive to uneven voxel or point densities. In contrast generating queries in the BEV domain is relatively simple. First, we generate a heatmap from $B_{out}^{’}$ with convolution. As outlined in [14], queries can be obtained by processing response scores using non-maximum suppression based on 3x3 max pooling, followed by the top-k selection. However, some hard targets, characterized by corrupted information such as sparse surface points or blurred image textures. After height compression, these features may be further deteriorated. Moreover, since targets are typically small in the BEV plane, their features may be overshadowed during convolution operations in the BEV backbone.
To address this, [48] propose a multi-stage query generation strategy. It groups queries into different difficulty levels and employs residual blocks to activate BEV features between levels to focus on harder instances. In this work, we simplify the approach by dividing queries into two groups: easy queries $Q_{e}$ and hard queries $Q_{h}$. The corresponding two-stage extraction process is detailed in following equations:
${H_{e}}=Heatmaphead({B_{out}})$ (8)
${Q_{e}}=Topk(NMS({B_{out}}))$ (9)
$M=I-Maxpooling(H)$ (10)
$H^{{}^{\prime}}=Heatmaphead(Resnet({B_{out}}))$ (11)
$Q_{h}=Topk(NMS(M\bullet{H^{{}^{\prime}}}))$ (12)
Where $M$ represents the mask generated from $Q_{e}$, which is used to obtain $Q_{h}$ by masking the regions corresponding to $Q_{e}$. When training, we compute the heatmap loss of the second stage only within the regions of $M$, guiding the network to focus on hard targets. However, activating BEV feature by the residual block has limited effectiveness—it can only aggregate local information from difficult regions to enhance responses. In the driving environment, the spatial relationships between objects exhibit certain regularities, such as the placement of vehicles and barriers. Based on this observation, we propose a hard instance activation (HIA) module, which leverages easy queries $Q_{e}$ to activate features in difficult regions. Specifically, we modify Equation 11 as follows:
${H^{{}^{\prime}}}=Heatmaphead(HIA({\color[rgb]{1,0,0}Q_{e}},{B_{out}}))$ (13)
As shown in Fig 6, in HIA, we first embed categories – in form of one-hot vectors – and positions into the easy queries. Then, we utilize self-attention to extract relationships between objects, resulting in enhanced queries $Q_{e}^{{}^{\prime}}$. Then, $B_{out}$ is flattened and regarded as query and $Q_{e}^{{}^{\prime}}$ serves as key and value. The query-to-BEV attention[49],[50] is utilized to propagate object relationships to every pixel in the BEV map. Subsequently, the results are reshaped back into a 2D tensor and passed through a residual block to obtain the refined BEV features $B_{out}^{{}^{\prime}}$. Finally, the easy and hard queries along with $B_{out}^{{}^{\prime}}$ are fed into the progressive decoder .
Figure 7: In the mixing block, queries are used to dynamically aggregate nearby voxel features from each modality. Subsequently, self-attention is employed to aggregate long-range voxel features among instances.
III-E Progressive Decoder
Many works [7, 50, 48, 9] only decode the BEV features at the end of their models. Although these features are rich in contextual and high-level semantic information, they lack fine-grained geometry-aware details, which limits the accuracy of box regression. To solve this, our progressive decoder (PQ) incorporates features from both BEV and voxel domains .
As shown in Fig 2, we decode the BEV features with several deformable attention blocks to obtain the refined queries $q^{{}^{\prime}}$. They are then fed into the (MMVFM) module to decode voxel features $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$. As depicted in Fig 7, two separate mixing modules are designed to dynamically aggregate voxel features. In each, we first use several linear layers to regress the proposal box $b_{i}$ from the query $q_{i}^{{}^{\prime}}$. Then, we divide the box into $G=g\times g\times g$ evenly spaced grid points $g_{1},…g_{G}$. Next, we extract grid features $Fg=[f_{g1},…,f_{gG}]$ from $V_{L}$ or $V_{I}$ using voxel pooling. We then embed offsets $\bigtriangleup g_{xk},\bigtriangleup g_{yk},\bigtriangleup g_{zk}$ of each grid point relative to the center of $b_{i}$ into $F_{g}$ through a linear layer, generating shape-aware grid features $F_{g}^{{}^{\prime}}\in\mathbb{R}^{G\times C}$. Following the Adamixer [51] operation, we sequentially perform channel mixing and spatial mixing on them, respectively:
$C_{i}=Linear(q_{i})$ (14)
$F_{g}^{{}^{\prime}}={C_{i}}\times F_{g}^{{}^{\prime}}$ (15)
$S_{i}=Linear(q_{i})$ (16)
$F_{g}^{{}^{\prime}}={S_{i}}\times transpose(F_{g}^{{}^{\prime}})$ (17)
Here, $C_{i}\in\mathbb{R}^{C\times C}$ and $S_{i}\in\mathbb{R}^{G\times G/4}$ represent the channel and spatial kernels, respectively. After the mixing operation, $F_{g}^{\prime}$ is flattened into a vector and passed through a linear layer to reduce the dimensionality to the final output $F_{I}$ or $F_{L}$. It is worth noting that $S_{i}$ differs in input and output channels from the original design of the adamixer $\mathbb{R}^{G\times G}$, which can reduce the number of parameters by compressing the spatial dimension. Following the aforementioned local mixing operation, we further aggregate long-range 3D features between objects with self-attention mechanisms. Finally, the features from both modalities, along with queries, are concatenated as the output. Subsequently, we decode the fused features through self-attention and a feed-forward layer. Finally, confidences and bounding boxes are regressed by a detection head.
IV EXPERIMENT
Our method is evaluated on the challenging NuScenes dataset. In this section, we first describe the dataset and implementation details, then report the evaluation performance of DDHFusion. Furthermore, we conduct ablation studies and comparison experiments with other homogeneous fusion methods for a thorough analysis.
IV-A Experimental Setup
Dataset. The nuScenes dataset is notable for its multi-sensor setup, which includes LiDAR, camera, radar, and GPS/IMU data, enabling comprehensive multi-modal research. It comprises 1,000 diverse driving scenarios, with a standardized split of 700 scenes for training, 150 for validation, and 150 for testing. It also features 10 object categories, such as cars, pedestrians, and cyclists, with over 1.4 million 3D bounding boxes annotated across various driving conditions. For 3D object detection, two key metrics are employed: mean Average Precision (mAP) and the nuScenes Detection Score (NDS), which collectively provide a holistic assessment of model performance. The mAP is computed by averaging the precision across 10 categories at distance thresholds of 0.5m, 1m, 2m, and 4m. NDS, on the other hand, is a composite metric that combines mAP with five True Positive (TP) error measures: mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), and mean Average Attribute Error (mAAE).
Implementation Details. Our model is implemented on the MMDetection3D framework. For the LiDAR backbone, we employ VoxelNet [52], while Swin-T serves as the image backbone. The detection range is set to [-54m, 54m] along x and y axes, and [-5m,3m] in the z direction. For ablation studies, the image size is set to $256\times 704$ for faster training. When evaluating the model on the validation or test sets, we follow previous works to set the image size to $384\times 1056$ and double the kernel channels in the LiDAR backbone to achieve better performance. For SAFS, we set score thresholds $d=0.01$, $s=0.25$ according to [41]. The max voxel number $N$ is set to $18000$. DDHFusion is trained on four 4090D GPUs with batch size of $12$. The training process is divided into two stages: (1) we first train the LiDAR-only network for 20 epochs. (2) Then, we combine the image and LiDAR branches and train the fused network for 6 epochs.
During training, we follow CBGS [51] to perform class-balanced sampling and employ the AdamW optimizer with max learning rate $2e^{-3}$. The copy-paste data augmentation [52] is applied to prevent overfitting. We also incorporate the multi-modal copy-paste augmentation[53] in the second stage to stabilize training process.
TABLE I: Evaluation results on the nuScenes test set. All models in the table are without ensemble or test-time augmentation. “L” denotes LiDAR and “C” denotes camera.
[…]
PointPillars [54] L 30.5 45.3 68.4 23.0 4.1 23.4 28.8 38.9 27.4 1.1 59.7 30.8
CenterPoint [55] L 60.3 67.3 85.2 53.5 20.0 56.0 63.6 71.1 59.5 30.7 84.6 78.4
TransFusion-L [14] L 65.5 70.2 86.2 86.6 28.2 50.8 58.8 60.0 32.9 9.5 77.5 88.4
LargeKernel3D [56] L 65.3 70.5 86.2 57.5 26.8 66.0 66.2 73.5 28.2 13.6 72.5 86.0
FocalFormer3D [48] L 68.7 72.6 87.8 59.4 37.8 65.7 73.0 77.8 77.4 52.4 90.0 83.4
PointPainting [1] LC 46.4 58.1 77.9 35.8 15.8 37.3 36.2 60.2 41.5 24.1 73.3 62.4
3D-CVF [15] LC 52.7 62.3 83.0 45.0 15.9 49.6 48.8 65.9 51.2 30.4 74.2 62.9
MVP [24] LC 66.4 70.5 86.8 58.5 26.1 57.3 67.4 74.8 70.0 49.3 89.1 85.0
AutoAlignV2 [57] LC 68.4 72.4 87.0 59.0 33.1 59.3 69.3 - 72.9 52.1 87.6 -
MSMDFusion [25] LC 71.5 74.0 88.4 61.0 35.2 66.2 71.4 80.7 76.9 58.3 90.6 88.1
TransFusion [14] LC 68.9 71.7 87.1 60.0 33.1 60.8 68.3 78.1 73.6 52.9 88.4 86.7
BEVFusion [7] LC 70.2 72.9 88.6 60.1 39.3 63.8 69.8 80.0 74.1 51.0 89.2 85.2
BEVFusion [8] LC 71.3 73.3 88.1 60.9 34.4 62.1 69.3 78.2 72.2 52.2 89.2 86.7
UVTR [12] LC 67.1 71.1 - - - - - - - - - -
GraphBEV [9] LC 71.7 73.6 89.2 60.0 40.8 72.1 64.5 80.1 76.8 53.3 90.9 88.9
FocalFormer3D [48] LC 71.6 73.9 88.5 61.4 35.9 66.4 71.7 79.3 80.3 57.1 89.7 85.3
CMT [58] LC 70.4 73.0 87.2 61.5 37.5 62.8 72.4 86.9 79.4 58.3 86.9 83.1
ObjectFusion [59] LC 71.0 73.3 89.4 59.0 40.5 71.8 63.1 76.6 78.1 53.2 90.7 87.7
DeepInteraction [57] LC 70.8 73.4 87.9 60.2 37.5 63.8 70.8 80.4 75.4 54.5 91.7 87.2
DeepInteraction++ [60] LC 72.0 74.4 - - - - - - - - - -
BiCo-Fusion [27] LC 72.4 74.5 88.1 61.9 38.2 73.3 65.7 80.4 78.9 59.8 89.7 88.3
DDHFusion LC
TABLE II: Evaluation results on the nuScenes val set. “L” denotes LiDAR and “C” denotes camera.
[…]
CenterPoint [55] L 60.3 67.3
TransFusion-L [14] L 65.5 70.2
TransFusion [14] LC 67.5 71.3
BEVFusion [7] LC 68.5 71.4
GraphBEV [9] LC 70.1 72.9
CMT [58] LC 67.9 70.8
MSMDFusion [25] LC 69.3 72.1
ObjectFusion [59] LC 69.8 72.3
UVTR [12] LC 65.4 70.2
DeepInteraction [57] LC 69.9 72.6
DeepInteraction++ [60] LC 70.6 73.3
BiCo-Fusion[27] LC 70.5 72.9
DDHFusion (ours) LC 71.6 73.8
DDHFusion-large (ours) LC 72.8 74.4
IV-B Results and Comparison
TABLE III: Ablation study of each module.
Methods HVF HBF PQG PD mAP NDS
(1) 68.4 71.2
(2) 70.0 72.1
(3) 69.1 71.8
(4) 70.6 72.6
(5) 71.4 73.1
(6) 71.6 73.8
TABLE IV: Ablation study of HVF.
Methods IV-Mamba CV-Mamba mAP NDS
(1) 68.2 71.2
(2) 69.4 71.9
(3) 69.6 71.8
(4) 70.0 72.3
TABLE V: Ablation study of different downsampling strides in HVF.
Method Voxel Stride mAP NDS
(1) [1] 69.6 71.9
(2) [1, 2] 69.8 72.2
(3) [1, 2, 4] 70.0 72.3
TABLE VI: Ablation study of HBF.
Methods IB-Mamba CB-Mamba mAP NDS
(1) 70.0 72.3
(2) 70.3 72.4
(3) 70.5 72.5
(4) 70.8 72.7
IV-C Ablation Studies
In this section, we demonstrate the effectiveness of each component through comprehensive ablation studies.
Ablation of each module As shown in Table V, we evaluate the contribution of each module to the detection performance. We replace the decoder in BEVFusion with deformable attention as the baseline. First, the proposed voxel fusion network significantly improves performance by 1.6 mAP and 0.9 NDS. Introducing only the BEV fusion network also provides a gain of 0.7 mAP and 0.5 NDS. As reported in the third row, the combination of both homogeneous fusion networks achieves a substantial improvement of 2.2 mAP and 1.4 NDS. Additionally, the progressive query generation module enhances object detection recall, contributing 0.7 mAP. Finally, integrating voxel decoder into the original BEV decoder to form the progressive decoder further increases performance by 0.3 mAP and 0.7 NDS. In summary, the four proposed components fully exploit the potential of dual-domain multi-modal features for high-performance object detection.
TABLE VII: Comparsion of different activation methods in PQG.
Activation Methods mAP NDS
baseline 70.8 72.7
ResNet[48] 71.0 72.8
HIA (ours) 71.4 73.0
TABLE VIII: Comparsion of different decoder numbers in MMVFM.
Methods BEV decoder N voxel decoder M mAP NDS
(1) 0 1 70.9 72.6
(2) 3 0 71.4 73.1
(3) 3 1 71.6 73.8
(4) 3 2 71.5 73.9
(5) 3 3 71.6 73.7
TABLE IX: Ablation study of MMVFM
Methods mixing block self-attention mAP NDS
(1) 71.4 73.0
(2) 71.5 73.3
(3) 71.6 73.4
(4) 71.6 73.8
TABLE X: Comparsion of image sizes and channels of LiDAR backbone.
Model Image Size or Channels mAP NDS
Image Backbone 256$\times$704
384$\times$1056
LiDAR Backbone (16,32,64,128)
(32,64,128,256)
Analysis of homogeneous voxel fusion network As shown in Table IV, we analyze the role of the Mamba-based modules in the voxel fusion stage. In the second row, we introduce a 3D U-Net with only IV-Mamba, without directly fusing image and LiDAR voxels. This approach improves performance by 1.2 mAP and 0.5 NDS, demonstrating that IV-Mamba can effectively capture long-range dependencies within each modality. The CV-Mamba can also enhance network performance, indicating the effectiveness of global fusion of voxels with different modalities and distributions. By combining both approaches with a simple merge-and-split strategy, we not only align multi-modal features but also preserve the relative independence of each modality, avoiding interference from noise or low-quality information from other modality, thereby further improving accuracy. As shown in Table V, setting the voxel strides to [1, 2, 4] in HVF achieves the best performance, as increasing the downsampling rate allows the network to capture structural information at different granularities.
Analysis of homogeneous BEV fusion network We further investigate HBF. As shown in Table VI, the SS2D module in the IB-Mamba enables full-scene feature propagation across the dense BEV map in four directions, resulting in improvements of 0.3 mAP and 0.1 NDS. Our proposed CB-Mamba achieves gains of 0.5 mAP and 0.2 NDS. It effectively propagates multi-modal features while adaptively exploring their latent relationships, thereby eliminating feature misalignment in the BEV domain. Notably, the combination of IB-Mamba and CB-Mamba yields the best performance, which benefits subsequent the query generation and decoder process.
Comparsion of different activation method In PQG, we explore different methods to activate BEV features for extracting harder queries in the second stage. As shown in Table VII, following the approach described in [48], utilizing a residual block yields a minor improvement of 0.2 mAP and 0.1 NDS. Our proposed HIA module further incorporates spatial cues from easy queries, enabling the identification of remaining hard objects in driving scenarios and achieving higher accuracy gains. However, it is important to note that the cross-attention between queries and BEV features introduces additional computational overhead, which limits the number of queries and stages. We plan to address this issue in the future work.
Analysis of progressive decoder In the progressive decoder, both BEV-based and voxel-based decoders are employed for query refinement. As shown in Table IX, we explore various configurations of their layer counts (M and N) to analyze their impact on performance. In the first row, using only a single voxel-based decoder yields suboptimal accuracy, even underperforming the second row, which adopts a 3-layer BEV decoder as in [48]. The BEV map at the end of the network contains richer contextual information and higher-level semantic features compared to voxels. Additionally, its dense representation is more robust to point cloud density variations than sparse voxel representation, both of which are critical for object detection in complex environments. However, as shown in the third row, adding a voxel decoder after the BEV decoder improves performance by 0.2 mAP and 0.7 NDS, indicating that incorporating voxel features with fine geometric details can enhance regression accuracy of predicted boxes. However, further increasing layer counts provides negligible performance gains. In Table X, we demonstrate that the mixing block and self-attention mechanism can adaptively aggregate shape features and long-range spatial information respectively, each contributing to performance improvements. As reported in the last row, combining them achieves the best results.
Analysis of on different scaling choices
IV-D Comparsion of Different Homogeneous Fusion Methods
We further compare DDHFusion with other three representative homogeneous fusion methods mentioned in the introduction, including GraphBEV, DeepInteraction++, and UVTR, through extensive experiments.
First, we evaluate their performance under different weather and lighting conditions. Since DeepInteraction++ and UVTR do not provide relevant data in their papers, we reproduce their models using open-source code to obtain the accuracy. As shown in Table XI, DDHFusion achieves the highest mAP across all environments. Unlike UVTR, DDHFusion is not negatively affected by rainy conditions and performs better even in poor lighting environments. This robustness stems from the comprehensive intra-modal and cross-modal feature learning in DDHFusion, which progressively filters out corrupt information to learn robust latent representations. Additionally, as shown in Table XII, we evaluate the performance of different models on object detection at varying distances, where “Near,” “Middle,” and “Far” represent ranges of $<20$m, [20[20[ 20m, $30$m], and $>30$m, respectively. Again, our proposed DDHFusion outperforms other methods across all distance ranges.
TABLE XI: Performance comparisons under different lighting and weather conditions
Method Sunny Rainy Day Night
GraphBEV[9] 70.1 70.2 69.7 45.1
Deepinteraction++[60] 70.3 70.4 70.6 42.9
UVTR [12] 66.4 65.1 65.5 36.7
DDHFusion (ours) 71.8 71.9 71.8 46.8
TABLE XII: Performance Comparisons with different ego distances
Method Near Middle Far
GraphBEV[9] 78.6 65.3 42.1
Deepinteraction++[60] 80.0 66.5 42.4
UVTR [12] 77.2 61.2 37.0
DDHFusion (ours) 82.3 67.8 44.2
TABLE XIII: Comparison of memory cost and inference speed.
Method Memory (MB) FPS
GraphBEV[9] 9851 5.7
Deepinteraction++[60] 13125 3.8
UVTR [12] 17640 3.6
DDHFusion (ours) 12189 4.2
Figure 8 showcases the detection results of DDHFusion, demonstrating its ability to accurately classify and localize objects of various categories in complex environments. Figure 9 provides a qualitative comparison with other models. As shown in the first row, DDHFusion successfully detects three distant and crowded objects. In the second row, bounding boxes generated by DDHFusion show minimal deviations. In the third row, it successfully identifies the small and distant object that other methods fail to detect. These results highlight the superiority of our dual-domain homogeneous fusion approach.
Figure 8: The visualization of DDHFusion on nuScenes validation set.
Figure 9: The visualization of different methods on nuScenes validation set.The green and blue boxes represent groundtruths and inference results respectively
Our method demonstrates competitive performance in terms of both memory consumption and runtime, highlighting its efficiency and practicality. For a fair comparison, all models are deployed in the mmdetection framework and tested on a single 4090D GPU. As shown in Table XIII, The computational resource consumption of DDHFusion is affordable, demonstrating the efficiency of our Mamba-based fusion modules. GraphBEV is the most lightweight, as it performs fusion only in the BEV domain. In contrast, UVTR incurs the highest computational burden due to its retention of dense voxels, while DDHFusion mitigates this issue through voxel sampling operations. Notably, DeepInteraction++ employs flash-attention for acceleration. Similarly, we believe that further hardware optimization of Mamba could significantly improve efficiency.
V Conclusion